OpenStudy (anonymous):

Which answer represents the domain of the logarithmic function given below? F(x) = log9 x A. x 0 B. x < 0 C. x > 0 D. All real numbers

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

you cannot take the log of a number that is not positive

OpenStudy (misty1212):

so find the one that says "positive numbers"

OpenStudy (misty1212):

\[\log_9(x)=y\iff 9^y=x\] and since \(9^y\) is always positive, that means \(x\) must always be positive

OpenStudy (anonymous):

a. is \[x \ge 0\]

OpenStudy (misty1212):

oh no

OpenStudy (misty1212):

you cannot take the log of zero since \(9^y>0\) always

OpenStudy (welshfella):

no 0

OpenStudy (anonymous):

so what is it

OpenStudy (danjs):

Graph y=log(x) and take a look what it does

OpenStudy (danjs):

notice the values x can take on

OpenStudy (danjs):



OpenStudy (danjs):

The valeues of the base just squishes it in the vertical Y direction ...